Kenny Harrison Corp., a capital goods manufacturing business that started Jan 4, 2008, and operates on a calendar- year basis. Plus, the use the installment sales method of profit recognition in accounting for all its sales. The following data were taken from 2008 and 2009 records. Item: 2008 2009 Installment Sales 480,000 620,000 Gross profit as a percent of costs 25% 28% Cash collections on sales of 2008 140,000 240,000 Cash collections on sales of 2009 0 180,000 Directions: A- Compute the amount of realized gross profit to be recognized on the 2009 income statement, using the installment sales method. B- Compute the amount of realized gross profit to be recognized on the income statement, using the cost-recovery method. Expert Answer
Answer #1

Installment sales of 2008 = 480,000
Installment sales of 2009 = 620,000

Gross profit % on cost 2008 = 25%
Gross profit % on cost 2009 = 28%

Calculation of Costs:
2008:
Let the costs be x.
Sales = Cost + Gross Profit
480,000 = x + 0.25x
480,000 = 1.25x
x = 480,000 / 1.25
x = 384,000 i.e. cost of goods sold in 2008.
Gross profit = 480,000 - 384,000 = 96,000

2009:
Let the cost be Y
Sales = Cost + Gross Profit
620,000 = Y + 0.28Y
620,000 = 1.28Y
Y = 620,000 / 1.28
Y = 484,375 i.e cost of goods sold
Gross Profit = 620,000 - 484,375 = 135,625

Gross Profit on sales = Gross profit / Sales
2008 : 96,000 / 480,000 = 0.2 or 20%
2009: 135,625 / 620,000 = 0.2188 or 21.88% rounded off.

A). Realised gross profit to be recognized on the 2009 income statement, using the installment sales method.
Realised profit = Cash Collection * Gross profit %
2009:
Realised profit = (240,000 * 20%) + (180,000 * 21.88%) = 87,384

B). Realized gross profit to be recognized on the income statement, using the cost-recovery method.
In 2008:
Total cost of goods sold = 384,000
Cash collected = 140,000
Balance cost to be recovered = 384,000 - 140,000 = 244,000
Gross profit = 0

In 2009:
2008 Cost of goods unrecovered = 244,000
Cash Collected for 2008 sales = 240,000
Balance unrecovered = 0
Gross profit of 2008 = 0

2009 Cost of Goods = 484,375
Cash Collected of 2009 sales = 180,000
Balance cost unrecovered = 484,375 - 180,000 = 304,375
Gross profit of 2009 sales = 0